OpenStudy (anonymous):

Draw a card 8 times from a shuffled deck, with replacement. What is the probability that at least 5 of the draws will be face cards?

OpenStudy (anonymous):

\[\left(\begin{matrix}n \\ k\end{matrix}\right)p^k(1-p)^{n-k}\]

OpenStudy (anonymous):

n=8 k=5 \[p =\frac{ 3 }{ 13 }\]

OpenStudy (anonymous):

somehow i got 0.01668 but the book's answer is 0.01941 did i mistake some numbers or I miss something inside the formula ?

OpenStudy (anonymous):

you have to repeat with \(P(x=6),P(x=7)\) and \(P(x=8)\) then add

OpenStudy (anonymous):

i would use wolfram

OpenStudy (kropot72):

You need to find the probability that exactly 6 will be face cards, then exactly 7 will be face cards and finally that exactly 8 will be face cards. Then add these 3 values of probability to the probability of exactly 5 cards thsat you have already found.

OpenStudy (anonymous):

k I will use wolfram would be long calculation

OpenStudy (anonymous):

with binomial [8,5](3/13)^5(1-3/13)^3+....

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

btw don't write \(1-\frac{3}{13}\) since that is clearly \(\frac{10}{13}\)

OpenStudy (anonymous):

i guess my calculation in wolfram is correct while this is a bit odd http://www.wolframalpha.com/input/?i=56%283%2F13%29^5%2810%2F13%29^3%2B28%283%2F13%29^6%2810%2F13%29^2%2B8%283%2F13%29^5%2810%2F13%29%2B%283%2F13%29^6

OpenStudy (anonymous):

thank you both for help @satellite73 and @kropot72

OpenStudy (anonymous):

closing this now because the question is already answer

OpenStudy (kropot72):

You're welcome :)